Previous | Calendar |
June |
2009 |
Next | ||||||||||||||||||||||||||||||
---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday | ||||||||||||||||||||||||||||
31JD151at ~13:30 MST, Data system crashes (Edgeport/8 dies) |
1JD152Site Visit (SpB, NT, SL, L, L, CG): Details , Photos (18) Span cylinder at 400 psi. N2 at 1080 psi. at ~10:00 MST, swapped out Edgeport/8, data system back up. Wet and drizzly (some thunder in distance). |
2JD153at 10:50 MST, the flow rate in the li-6262 system drops significantly (check the GAST pump and/or filters)... |
3JD154 |
4JD155Site Visit (SpB, LL, SL, LY): Details , Photos (9) Span cylinder at 300 psi. N2 at 900 psi. from 11:00-11:26 MST, replaced 2-micron Nupro filters in LI-6262 system (both at inlet and in the box) PAR installation work. Warm and pleasant (some clouds in the afternoon and a light sprinkle of rain). |
5JD156 |
6JD157at 17:20 MST, chan 201 (fast logger) stops sending out data
|
7JD158Site Visit (SpB): Details , Photos (?) at ~11:30 MST, replaced 12V batteries in "fast" cr23x (fixes chan 201). Big clouds around; some snow (frozen pellets), gusty and variable winds.
|
8JD159at 15:00 MST, large lightning strike, data system goes down
|
9JD160Site Visit (SpB, LL, LY, RW, RD, +): Details , Photos (26) Span cylinder at 250 psi. N2 at 705 psi. at ~9:30 MST, swapped out Edgeport/8, data system back up (note, left channels 200 & 202 off system). And replaced short-haul for UC logger. Partly cloudy. Rain/storm never materialized at the site. Huge Cb on plains.
|
10JD161
|
11JD162
|
12JD163
|
13JD164
|
14JD165
|
15JD166
|
16JD167
|
17JD168Site Visit (SpB, NT, SL, LY, RW, LL, CG, +): Details , Photos (17) Span cylinder at 2000 psi. N2 at 510 psi. at ~10:30 MST, added channels 200 (UC logger) & 202 (ucb logger) back to the data system. at ~11:15 MST, swapped span gas (old cylinder=SG9151924 at 150 psi, new cylinder= CC245057, at 2000 psi) from 12:25-12:30 MST, swiped krypton hygrometer Warm in the morning. Some clouding up in the afternoon, but no rain (until after we left).
|
18JD169
|
19JD170
|
20JD171
|
21JD172Summer Solstice
|
22JD173
|
23JD174Note, dew on 2.5m csat in morning until around 10:00 MDT. (others seem fine, no rain)
|
24JD175
|
25JD176Site Visit (SpB, SL, LL, R, +): Details , Photos (11) Span cylinder at 1800 psi. N2 at 300 psi. from 10-10:50 MST, co2b.ec is now h2o.ec; made a drastic change to h2o.ec offset using dial; also changed range of h2o.ec to 5000 mV, now running fast_08.csi (see details) at ~10:30 MST, removed channel 200 (UC logger) from the data system. Fairly calm. Dark clouds were all around. It rained briefly at around 11. Then we left. Gray Jeep returned for repairs; replaced with Suburban
|
26JD177Report Jeep Mileage to Lease Operations
|
27JD178
|
28JD179
|
29JD180Site Visit (SpB, CG, AJ, SO, L): Details , Photos (25) Span cylinder at 1800 psi. N2 at 200 psi --> 2000 psi. at ~10:30 MST, reconnected channel 200 (UC logger); added in-line surge protector at tower end. from 13:00-14:10 MST, switched N2 cylinder at the tower (new cylinder=CC199722 (2000 psi); old cylinder=CC198519 (200 psi)) from 12:00-12:30 MST, made more adjustments to h2o.ec offset using dial (see details) from 13:20-13:45 MST, replaced domes on REBS Rnet sensor TRAM was running! Hot and Dry. Cu clouds formed as the day went on. It rained briefly at around 13:30 MST.
|
30JD181
|
1JD182
|
2JD183
|
3JD184
|
4JD185 |